oregonduck (oregonduck):

What will be the perimeter and the area of the rectangle below if it is enlarged using a scale factor of 5.5? A rectangle is shown. The length of the rectangle is labeled as 8 cm, and the width is labeled as 6 cm. Perimeter = 50 cm, area = 155.25 cm2 Perimeter = 50 cm, area = 1,452 cm2 Perimeter = 154 cm, area = 155.25 cm2 Perimeter = 154 cm, area = 1,452 cm2

oregonduck (oregonduck):

I am pretty sure it is d

oregonduck (oregonduck):

Because 8*5.5=44 and 6*5.5=33 then 44+44+33+33= 154 cm And then 44*33=1,452 cm^2

OpenStudy (anonymous):

ok thats right now solve for perimeter

OpenStudy (anonymous):

nvm u did and your right

oregonduck (oregonduck):

i did it is the first part

OpenStudy (just_one_last_goodbye):

I think thats right @OregonDuck Perimeter = 154 cm, area = 1,452 cm2

OpenStudy (anonymous):

it is right

OpenStudy (just_one_last_goodbye):

When I do the math thats what I get ^_^

oregonduck (oregonduck):

yup, ok thxs guys i have some more, yeah same here @just_one_last_goodbye